OpenStudy (anonymous):

Consider the piecewise continuous function $\left\{\begin{matrix} 2t-5 &0\leq t<3 \\ e^{-2t}& 3\leq t<5\\ t^2+3t&5\leq t<6 \\ -2&t\geq 6 \end{matrix}\right.$ Write f(t) in terms of unit step functions and find the laplace transform of f(t)

OpenStudy (ahsome):

Instead of \($\), use \[ instead. So I can see what you are talking about

OpenStudy (anonymous):

\[\begin{matrix} 2t-5 &0\leq t<3 \\ e^{-2t}& 3\leq t<5\\ t^2+3t&5\leq t<6 \\ -2&t\geq 6 \end{matrix}\]

OpenStudy (anonymous):

@satellite73 Any help?